# Utility Boundary of Dataset Distillation: Scaling and Configuration-Coverage Laws

### v1

@@ C.4 Proof of Key Insights
(ii) The remaining terms vanish at canonical rates: $L_{R}\rho_{a}^{T}\|\delta_{0}\|=O(T^{-\beta})$ under PL (karimi2016linear); $e_{g}=O(1/\sqrt{k}+1/\sqrt{n})$ by Lemma C.4; and $e_{\mathrm{te}}=O(1/\sqrt{m})$ by Lemma C.5.
@@ Appendix D Proofs for the Coverage–Aware Bounds
where $c_{1},c_{2},c_{3}>0$ are universal constants (depending only on the choice of empirical Bernstein inequality; see, e.g., boucheron2013concentration, theorem 2.10).
@@
By symmetrization (e.g. mohri2018foundations, theorem 3.1),
@@
We refine the concentration step by replacing Hoeffding/McDiarmid with an empirical-Bernstein deviation for bounded variables. For a fixed $f$, boucheron2013concentration, theorem 2.10 implies that for any $\delta\in(0,1)$,
@@
To make Eq. (31) uniform over $f\in\mathcal{F}_{i}$, we proceed by localization via symmetrization: for any $r>0$, define the localized class $\mathcal{F}_{i}(r):=\{f\in\mathcal{F}_{i}:\mathrm{Var}(f(Z))\leq r\}.$ By the same symmetrization step as in Eq. (30), applied to the truncated excess loss $f-\mathbb{E}f$ and then peeling over dyadic radii $r_{m}=2^{-m}B^{2}$, we obtain (see, e.g., mohri2018foundations, section 3.5) that with probability at least $1-\delta$,
@@
If the distilled set $\mathcal{D}_{s}$ is generated from (or depends on) $\mathcal{D}_{\tau}$, the Hoeffding-type bound used for the distilled-side sampling deviation in Eq. (39) should be replaced by a high-probability information-theoretic tail. By bu2020tightening (see also xu2017information; steinke2020reasoning), if the class is bounded by $B_{g}$ (thus sub-Gaussian with proxy $B_{g}$), there exists a universal constant $C_{I}^{\prime}>0$ such that, with probability at least $1-\varepsilon/2$,
@@
(High-probability variant). If the distilled dataset $\mathcal{D}_{s}$ can depend on the real dataset $\mathcal{D}_{\tau}$, the distilled-side sampling deviation in Eq. (60) should be replaced by a high-probability information-theoretic tail (e.g., bu2020tightening, theorem 7; cf. xu2017information; steinke2020reasoning). There exists a universal constant $C_{I}^{\prime}>0$ such that, with probability at least $1-\varepsilon$,
@@
A high-probability mutual-information tail (e.g., bu2020tightening; xu2017information; steinke2020reasoning) implies that for any fixed $a$ and any $\varepsilon\in(0,1)$,
@@ E.1 Setup and recall assumptions
This is the standard descent lemma for $L$-smooth functions (nesterov2013introductory; bubeck2015convex).

### v3

@@ C.4 Proof of Key Insights
(ii) The remaining terms vanish at canonical rates: $L_{R}\rho_{a}^{T}\|\delta_{0}\|=O(T^{-\beta})$ under PL (Karimi et al. 2016); $e_{g}=O(1/\sqrt{k}+1/\sqrt{n})$ by Lemma C.4; and $e_{\mathrm{te}}=O(1/\sqrt{m})$ by Lemma C.5.
@@ Appendix D Proofs for the Coverage–Aware Bounds
where $c_{1},c_{2},c_{3}>0$ are universal constants (depending only on the choice of empirical Bernstein inequality; see, e.g., Boucheron et al., 2013, theorem 2.10).
@@
By symmetrization (e.g. Mohri et al., 2018, theorem 3.1),
@@
We refine the concentration step by replacing Hoeffding/McDiarmid with an empirical-Bernstein deviation for bounded variables. For a fixed $f$, Boucheron et al., 2013, theorem 2.10 implies that for any $\delta\in(0,1)$,
@@
To make Eq. (31) uniform over $f\in\mathcal{F}_{i}$, we proceed by localization via symmetrization: for any $r>0$, define the localized class $\mathcal{F}_{i}(r):=\{f\in\mathcal{F}_{i}:\mathrm{Var}(f(Z))\leq r\}.$ By the same symmetrization step as in Eq. (30), applied to the truncated excess loss $f-\mathbb{E}f$ and then peeling over dyadic radii $r_{m}=2^{-m}B^{2}$, we obtain (see, e.g., Mohri et al., 2018, section 3.5) that with probability at least $1-\delta$,
@@
If the distilled set $\mathcal{D}_{s}$ is generated from (or depends on) $\mathcal{D}_{\tau}$, the Hoeffding-type bound used for the distilled-side sampling deviation in Eq. (39) should be replaced by a high-probability information-theoretic tail. By Bu et al. (2020, theorem 7) (see also Xu & Raginsky, 2017; Steinke & Zakynthinou, 2020), if the class is bounded by $B_{g}$ (thus sub-Gaussian with proxy $B_{g}$), there exists a universal constant $C_{I}^{\prime}>0$ such that, with probability at least $1-\varepsilon/2$,
@@
(High-probability variant). If the distilled dataset $\mathcal{D}_{s}$ can depend on the real dataset $\mathcal{D}_{\tau}$, the distilled-side sampling deviation in Eq. (60) should be replaced by a high-probability information-theoretic tail (e.g., Bu et al., 2020, theorem 7; cf. Xu & Raginsky, 2017; Steinke & Zakynthinou, 2020). There exists a universal constant $C_{I}^{\prime}>0$ such that, with probability at least $1-\varepsilon$,
@@
A high-probability mutual-information tail (e.g., Bu et al. 2020; Xu & Raginsky 2017; Steinke & Zakynthinou 2020) implies that for any fixed $a$ and any $\varepsilon\in(0,1)$,
@@ E.1 Setup and recall assumptions
This is the standard descent lemma for $L$-smooth functions (Nesterov, 2013; Bubeck et al., 2015).